The half life of a radioactive substance is 30 days and it takes 60 days for disintegrating to ¾ th of its original mass what is its average life

Grade:10

1 Answers

Neeti
8 years ago
we know that T1/2 = 0.693/k
We have T1/2 so we get K
and Tavg = 1/k so putting the value of k we get Tavg.
:)